# Synthetic Geology - Structural Geology Meets Deep Learning

### v1

@@ Methods
Figure 1: An example of geological history illustrating the application of transformations and depositions. (A) The initial strata ${\bf m}_{0}({\bf x})$, (B) folded by a transformation $\mathcal{T}_{\text{fold}}$, (C) intruded by a dike $\mathcal{D}_{\text{dike}}$, and (D) finally faulted by another transformation $\mathcal{T}_{\text{fault}}$. Each arrow represents a process $\mathcal{P}_{i}$ in the overall history $\mathcal{H}$. In practice, these processes are sequenced randomly from a set ${\cal P}_{j},j=1,\ldots,k$ using a Markov chain, leading to varied geological outcomes.
Figure 2: Illustration of the pipeline for random geological model generation. A batch request triggers a Markov chain sampler, which selects and sequences geological processes (e.g., folding, deposition) from a transition matrix. For each process, parameters such as amplitude and wavelength are drawn from tuned random variables. Finally, the constructed computational sequences are applied to produce multiple models in the form of 3D tensors.
The above simulation can generate a virtually infinite number of geological models with variable complexity and at any specified resolution, which can be used as data for a generative model. Nonetheless, just like any other simulation, it depends on a number of hyperparameters that need to be carefully tuned. Every geological process depends on a set of parameters that are sampled as random variables and must be tuned such that the outcome of each geological process is "realistic" as visually determined by geologists. In addition, the transition probability $\pi_{j}$ of every geological process must be chosen such that it yields a reasonable geology as judged by a domain expert. However, it is important to realize that such models represent what experts consider to be unbiased geology. Since there are no direct observations on the geology at the (3D) resolution and scale that we need, it is virtually impossible to know if such models introduce implicit bias. Nonetheless, such models are extremely useful in generating what experts believe is a realistic Earth model.
Synthetic Geology Using Generative AI
Let ${\bf m}({\bf x})$ be a geological model generated by the simulation process using a sampled geological history ${\cal H}({\bf m}_{0})$. These models reside in a high-dimensional probability space $\pi_{m}$ covering the range of all possible 3D geological structures.
Given a method to generate virtually infinite samples of Geomodels, ${\bf m}({\bf x})$, we now have the required tools to train a foundation model to learn the probability of any given Geomodel, $\pi_{{\bf m}}({\bf m})$ or its associated score function defined as, $s_{m}=-{\bm{\nabla}\,}_{m}\log(\pi_{m}({\bf m}))$.
While the original simulation process can generate diverse Geomodels, it is hard, and in some cases virtually impossible, to use the simulation for any other application besides the forward process of generating a random 3D geological volume from a parameterized history ${\cal H}({\bf m}_{0})$. In contrast, foundation models support a broader range of applications that vary from generation to conditional generation or inversion, enabling the incorporation of other data or constraints based on conditional distributions [27].
For instance, while the simulation process requires a structured history of geological events, a foundation model can generate realistic samples conditioned directly on partial observations in the spatial domain ${\bf x}$, such as borehole data or surface measurements. Since our ultimate goal is to fuse multiple sources of geo-scientific data, we seek to replace the rigid simulation process with a flexible, trainable foundation model.
One approach to learning a probability density function $\pi_{\bf m}$ is the denoising or diffusion model, now commonly used for image and movie generation. While diffusion methods are often applied to continuous-valued data, the generation of discrete-valued data has only more recently been addressed [28] through the theory first proposed by Hyvärinen [29]. One of the difficulties with score matching for our application is the fact that the score does not exist for discrete distributions. A more general way to approach the problem of matching distributions that does not require the score can be derived through flow matching [10].
Embedding Discrete Categorical Data
@@
To adapt flow matching to discrete categorical data, we reformulate the problem by embedding the data into a continuous space without imposing any inherent ordering on the categories. Instead of representing a geological model by ${\bf m}:R^{3}\rightarrow{\mathbb{N}}$ we introduce an encoded vector function
${\bf m}^{(oh)}:R^{3}\rightarrow[0,1]^{\mathbb{N}},$ (6)
such that each integer value of ${\bf m}$ is mapped to a new vector. The new function, ${\bf m}^{(oh)}$, represents the likelihood of a point ${\bf x}$ in our model to belong to a particular class that ranges from $1$ to ${\mathbb{N}}$. The vector function ${\bf m}_{i}^{(oh)}({\bf x})$ is sometimes referred to as the one-hot encoding of the model ${\bf m}$.
We further normalize these embedding vectors to have an unweighted mean of zero and unit length such that
${\bf m}_{i}^{(e)}=\frac{{\bf m}_{i}^{(oh)}-\left(\frac{1}{N},...,\frac{1}{N}\right)}{\|{\bf m}_{i}^{(oh)}-\left(\frac{1}{N},...,\frac{1}{N}\right)\|}.$ (7)
@@
Because we no longer require each embedding vector to remain on the probability simplex, the flow field is free to traverse an unconstrained continuous space. It is then the task of the model to learn the probability distribution $\pi_{m^{(e)}}$ in the embedded space, from which we can decode back to the original categorical space of ${\bf m}$. To decode a voxel $x^{(e)}$ back to a discrete category, we select the embedding vector ${\bf m}_{i}^{(e)}$ with the largest cosine similarity:
$i=\text{arg}\max_{i}\langle x^{(e)},{\bf m}_{i}^{(e)}\rangle.$ (8)
Flow Matching
Flow matching is a powerful technique that can be used to connect any two probability spaces. We define the distributions, ${\bf m}_{t}^{(e)}$ by
${\bf m}_{t}^{(e)}({\bf x})=t{\bf m}^{(e)}({\bf x})+(1-t){\bf z},$ (9)
where ${\bf z}\sim N(0,{\bf I})$. Flow matching learns the velocity field ${\bf v}=d/dt\left({\bf m}_{t}^{(e)}({\bf x})\right)={\bf m}^{(e)}({\bf x})-{\bf z}$ characterized by the unique minimizer of the quadratic objective
$\min_{{\bm{\xi}}}{\frac{1}{2}}{\mathbb{E}}_{t,{\bf z}}\,\left\|{\bf v}({\bf m}_{t}^{(e)},t;{\bm{\xi}})-{\bf m}^{(e)}({\bf x})+{\bf z}\right\|^{2}.$ (10)
Here ${\bf v}({\bf m}^{e}_{t},t;{\bm{\xi}})$ is a neural network model that takes ${\bf m}_{t}^{(e)}$ and $t$ as inputs and parameterized by ${\bm{\xi}}$. The parameters ${\bm{\xi}}$ are adjusted to solve the optimization problem, and this is done through stochastic gradient descent.
While in standard training of generative models, the data set is finite and static, our data set is practically infinite and dynamic. This stems from the ability to generate models "on the fly" as the system is being generated, allowing for a diverse set of models to train the system.
Applications
@@
A trained model ${\bf v}({\bf m}_{t},t;{\bm{\xi}})$ indirectly learns the distribution $\pi_{\bf m}$, allowing us to draw or infer new geological models that are similar to the training set. Generating arbitrary samples from $\pi_{\bf m}$ is an example of unconditional generation where the flow process is not guided or conditioned by any constraints. To draw a new sample, we integrate the learned ODE using a numerical solver such as Runge-Kutta or adaptive Heun’s method,
${\frac{d{\bf m}}{dt}}={\bf v}({\bf m}_{t},t;{\bm{\xi}})\quad{\bf m}(0)={\bf m}_{0}\sim N(0,{\bf I})\ \quad t\in[0,1].$ (11)
While the unconditional generation of models is helpful in many applications where models are needed for the simulation of other data (see, e.g., [30]), a more important application is the integration of measured data. Consider some noisy data of the form
@@
where ${\bf A}$ is the so called forward mapping and ${\bm{\epsilon}}\sim N(0,\sigma^{2}{\bf I})$ is the noise.
@@
There are a number of ways to integrate the measured data into the ODE system, Note that equation equation 12 is an algebraic constraint that is added to the ordinary differential equation. This changes the ODE into a Differential Algebraic Equation (DAE) [31]. The incorporation of such constraints in the context of deep learning was recently discussed in [32]. A straightforward approach is adding a term to the ODE that "attracts" the solution of the ODE toward the data, obtaining the system
${\frac{d{\bf m}}{dt}}={\bf v}({\bf m}_{t},t;{\bm{\xi}})-\mu{\bf A}^{\top}({\bf A}{\bf m}-{\bf d})\quad{\bf m}(0)={\bf m}_{0}\sim N(0,{\bf I})\ \quad t\in[0,1].$ (13)
The parameter $\mu$ balances between data fit and fit to the prior model and is tuned typically by cross-validation. For a general application of conditional data to the generation process, we add an additional embedding of conditional data ${\bf d}$ to the machine learning model and a modification to the loss function using a method adapted from recent work by Ahamed and Haber [33]. Ultimately, the goal is to learn to sample from the distribution $\pi_{\bf m}({\bf m}|{\bf d})$ to produce many different candidate geological models for a set of field data.
Results
Unconditional Generation of Earth Models
@@ Results
In our first set of experiments, we show that a flow matching model trained with suitable synthetic data in a continuously embedded space can generate realistic Earth models. Unconditional generation is carried out using the learned velocity flow field and a numerical ODE solver [34]. After integration, the resulting sample is decoded into the categorical space. Figure 3 highlights three example models produced with this approach.
@@
Figure 3: Three examples of unconditional Earth models. Panels (a–c) $64\times 64\times 64$ resolution generative samples using flow matching. Each model represents a $3.84\,\text{km}\,\times 3.84\,\text{km}\,\times 3.84\,\text{km}$ volume of the Earth’s crust.
@@
These models are diverse and include different topography, layer thickness, layer geometry, as well as dykes and intrusions. Such samples can support data generation [30] for various applications as discussed in [35, 36] as well as for training other machine learning tasks such as geological segmentation and mapping [37], or classification [38].
Conditional Generation of Earth Models
@@
The standard technique for solving this problem is kriging [39] and, more recently, interval kriging [40], which estimates lithological probabilities between boreholes by accounting for vertical and horizontal variability in rock types. However, kriging can struggle when the data size is small (few boreholes) and has the tendency to smooth interfaces when applied to categorical data. We demonstrate an alternative reconstruction method using a conditional flow matching model to interpolate between sparse boreholes. Because the technique is inherently generative, we can produce many different models that are consistent with the observed data and also with the underlying probability density function of the model.
@@
(a) t = 0.
(b) t = 0.4.
(c) t = 0.6.
(d) t = 0.8.
(e) t = 1.
Figure 5: A sequential visualization of the model shows the generative process that maps from Gaussian noise to plausible geology conditioned on sparse data. Decoded models in the discrete categorical space from intermediary distributions between $t=0$ to $t=1$ show snapshots from a continuous flow bridging between a Gaussian distribution and $\pi_{\bf m}({\bf m}|{\bf d})$.
@@
Figure 6: Three geological models from $\pi_{\bf m}({\bf m}|{\bf d})$ via numerical integration conditioned on available lithological data. The models offer a complete 3D $(64\times 64\times 64)$ resolution reconstruction, with the exterior surface of the models shown here for comparison.
@@
Conditional Feature Estimation and Reconstruction
Individual features of a reconstructed model, such as dike formations, can be analyzed in greater detail using an ensemble of samples from the conditional distribution $\pi_{\bf m}({\bf m})$. The planar dikes in the original model, the sparse dike data, and several reconstructions of their form are shown in Figure 7. The realizations show a high degree of variability due to the very limited conditional data.
(a) Dike material in true model and conditional data.
(b) 12 different dike realizations conditioned on surface and borehole data.
Figure 7: Comparison of conditioned dike realizations. In (a), the true shape of two geological dikes is shown, along with the sparse data sampled from surface and borehole data. (b) Many conditional realizations of the subsurface dike distribution are estimated by sampling from the generative model with the conditional data.
A large ensemble of samples can be used as a probability estimate for the likelihood of any individual voxel being part of the larger dike structure. A pattern emerges when plotting the extent of the planar dike based on the contour shape of probabilistic confidence intervals. Shown in Figure 8, a lower confidence interval corresponds to a larger and more aggressive extrapolation range from the existing boreholes. The smaller dike shown in the model has insufficient conditional data to produce an accurate reconstruction, a problem exacerbated by the true extent of the dike being largely obscured by sediment.
(a) Dike 1 probability map, isometric.
(b) Dike 2 probability map, isometric.
(c) Dike 1 probability map, front view.
(d) Dike 2 probability map, front view.

### v3

@@ 2.3 Synthetic Geology Using Generative AI
2.3.1 Flow Matching
Flow matching, also know as stochastic interpolation, is a powerful technique to learn a continuous normalizing flow between two probability distributions [lipman2023flow, albergoStochasticInterpolantsUnifying2023]. Rather than attempt to directly estimate a complex probability distribution such as $\pi_{\bf m}$, the model learns a velocity field that transports samples from a simpler source distribution $\pi_{0}$ at $t=0$ to our target distribution $\pi_{\bf m}$ at t=1t\!=\!1.
Let $t\in[0,1]$ be the time parameter, let ${\bf m}^{(e)}=\text{enc}({\bf m}({\bf x}))$ denote the continuous latent embedding of ${\bf m}({\bf x})\sim\pi_{\bf m}$ (see Equation (11)), and let ${\bf z}\sim N(0,{\bf I})$ be a Gaussian random variable drawn from the source distribution. Then the stochastic interpolant ${\bf m}_{t}^{(e)}$ between the start and target distributions is defined as
${\bf m}_{t}^{(e)}=t{\bf m}^{(e)}+(1-t){\bf z},$ (6)
and the velocity between the source and target is
${\bf v}_{t}=\frac{d}{dt}{\bf m}_{t}^{(e)}={\bf m}^{(e)}-{\bf z}.$ (7)
Flow matching seeks to learn the unique minimizer over all $t,\mathbf{z},\mathbf{m}^{(e)}$ of the quadratic objective
$\min_{\bm{\xi}}\frac{1}{2}\,\mathbb{E}_{t,\mathbf{z},\mathbf{m}^{(e)}}\left\|\mathbf{v}(\mathbf{m}_{t}^{(e)},t;\bm{\xi})-\mathbf{v}_{t}\right\|_{2}^{2}.$ (8)
In practice the true minimizer is approximated by a deep neural network
$\hat{\mathbf{v}}(\mathbf{m}_{t}^{(e)},t;\bm{\xi}),$ (9)
parameterized by $\bm{\xi}$ and trained using stochastic gradient descent.
At inference time, samples are drawn from $\pi_{0}$ and integrated through the learned velocity field $\hat{\bf v}$ using either an ordinary differential equation (ODE) or stochastic differential equation (SDE) solver to get novel samples from the embedded space of $\pi_{\bf m}$. In our method we use the deterministic ODE method with $\pi_{0}=N(0,{\bf I})$, as it is easier to work with and does not appear to compromise sample quality.
2.3.2 Embedding Discrete Categorical Data
@@
To adapt flow matching to discrete categorical geology models, we embed the model into a continuous vector space without imposing any ordering on the $N$ categories. A Geomodel is a discrete field ${\bf m}:\mathbb{R}^{3}\rightarrow\{1,\ldots,N\},$ assigning to each location $\bf x$} a categorical lithology label.
We map each category to a normalized embedding vector through
$\text{enc}(i)=\frac{e_{i}-\frac{1}{N}\mathbf{1}}{1-\frac{1}{N}},$ (10)
where $e_{i}$ is the $i$-th standard basis vector and $\mathbf{1}$ is the length-$N$ vector of ones. The embedded model is obtained pointwise as
${\bf m}^{(e)}({\bf x})=\text{enc}({\bf m}({\bf x})).$ (11)
Applying enc(.) across the entire domain yields a continuous field ${\bf m}^{(e)}\in\mathbb{R}^{[N,X,Y,Z]}$ on a mesh of size $[X,Y,Z]$. This continuous representation ${\bf m}^{(e)}$ resides in a highdimensional space for which we can compute gradients with respect to the embedded model parameters ${\bf m}^{(e)}$ for flow matching, see Equation (8).
@@
Because we do not constrain the embedded vectors to the probability simplex, the flow field is free to traverse an unconstrained continuous space. It is then the task of the model to learn the equivalent probability distribution $\pi_{{\bf m}^{(e)}}$ in the embedded space, from which we can decode back to the original categorical space of $\bf m$}.
To decode a voxel ${\bf m}^{(e)}({\bf x})\in\mathbb{R}^{N}$ back to a discrete category ${\bf m}({\bf x})=i\in\mathbb{N}$, we select the embedding vector over all $j\in\mathbb{N}$ with the largest cosine similarity:
$\text{dec}({\bf m}^{(e)}({\bf x}))=\arg\max_{j}\langle{\bf m}^{(e)}({\bf x}),\text{enc}(j)\rangle.$ (12)
Since the embedding vectors are unit-norm, this is equivalent to nearest-neighbor or shortest euclidean distance.
2.4 Applications to Inverse Problems
@@
A trained model $\hat{\bf v}({\bf m}_{t}^{(e)},t;{\bm{\xi}})$ indirectly learns the distribution $\pi_{\bf m}$ through the latent embedding space, allowing us to draw or infer new geological models that are similar to the training set. Generating arbitrary samples from $\pi_{\bf m}$ is an example of unconditional generation where the flow process is not guided or conditioned by any constraints. To draw a new sample, we integrate the learned ordinary differential equation (ODE)
${\frac{d{\bf m}}{dt}}=\hat{\bf v}({\bf m}_{t}^{(e)},t;{\bm{\xi}})\quad{\bf m}_{0}^{(e)}\sim N(0,{\bf I})\ \quad t\in[0,1],$ (13)
starting from pure Gaussian noise at $t=0$, using a numerical solver such as Runge-Kutta or adaptive Heun’s method.
[…]
@@
where $\bf A$} is the forward mapping [R2-C4-2], ${\bm{\epsilon}}\sim N(0,\sigma^{2}{\bf I})$ is Gaussian noise, and $\bf d$} is an indirect observation of $\bf m$}.
@@
Equation (14) is an algebraic constraint that is added to the original ODE from Equation (13), turning it into a Differential Algebraic Equation (DAE) [AscherPetzoldODEs], a formulation that recently has been explored in deep learning contexts [Boesen2025DAEs]. A straightforward approach is adding a term to the ODE that attracts the solution towards the measured data, obtaining the system
${\frac{d{\bf m}}{dt}}={\bf v}({\bf m}_{t},t;{\bm{\xi}})-\mu{\bf A}^{\top}({\bf A}{\bf m}-{\bf d})\quad{\bf m}_{0}\sim N(0,{\bf I})\ \quad t\in[0,1].$ (15)
The parameter $\mu$ balances between data fit and fit to the prior model and is tuned typically by cross-validation.
2.4.1 Conditional Flow Matching
[…]
$\hat{\bf d}={\bf A}\hat{\bf m}={\bf A}\left({\bf m}_{t}+(1-t)\cdot\hat{\bf v}({\bf m}_{t},{\bf d},t;{\bm{\xi}})\right)$ (16)
provides a differentiable estimate of the observable response. Since this has high variance for $t\rightarrow 0$ but must match the data as $t\rightarrow 1$, we introduce a time-weighted conditional loss,
$\mathcal{L}_{\text{cond}}=\left\|\hat{\bf v}({\bf m}_{t},{\bf d},t;{\bm{\xi}})-{\bf v}_{t}\right\|_{2}^{2}+\lambda\cdot t\cdot\left\|\hat{\bf d}-{\bf d}\right\|_{2}^{2}.$ (17)
The multiplying factor of $t$ ensures that the data consistency is emphasized only near the end of the flow trajectory, preventing instabilities in training. This extends CFM to inverse problems by jointly learning both the conditional velocity field and the reconstruction implicit in $\hat{\bf d}$.
Ultimately, the goal is to learn to sample from the posterior distribution $\pi_{{\bf m}|{\bf d}}$ to produce many different candidate geological models consistent in $\bf d$}.
2.5 Network Architecture
So far we have presented a method of reconstructing samples from both $\pi_{\bf m}$ and $\pi_{{\bf m}|{\bf d}}$. Now we describe our implementation of a deep learning network for modeling the conditional estimator $\hat{\bf v}({\bf m}_{t}^{(e)},{\bf d},t;{\bm{\xi}})$, noting that the unconditional generation operates with a similar architecture but without conditional embeddings[R2-C5-1]. To learn $\hat{\bf v}$, we use a magnitude preserving attention 3D U-Net [vaswaniAttentionAllYou2017, ronneberger2015u, Karras2023AnalyzingAI]. The model combines 3D convolutional skip connections and attention layers to capture multiscale relationships, both locally and at long range. Self-attention layers are particularly beneficial for geological constructions where long-range stratigraphic continuity and spatial correlation cannot be captured solely by convolutional operations.
The model requires a scheme for embedding and injecting both ${\bf A}^{\top}{\bf d}$ and $t$. At each resolution scale of the U-Net, a module [R1-C2] composed of convolution and activation layers provides feature extraction, and rescaling of $({\bf A}^{\top}{\bf d})^{(e)}$. We perform time embedding with either sinusoidal [vaswaniAttentionAllYou2017], random Fourier [rahimirandfour], or learnable Fourier features [tancikFourierFeaturesLet2020] which are then passed through a multi-layer perceptron (MLP). Finally a mixing module concatenates feature tensor ${\bf x}_{n}$ with the embedded adjoint ${\bf A}^{\top}{\bf d}$, then applies a feature-wise linear modulation (FiLM) through a residual connection [perez2018film]
$\mathbf{x}_{n+1}=\mathbf{x}_{n}+\mathcal{F}\left(\underbrace{(\mathbf{1}+\bm{\gamma}_{t})}_{\text{time scale}}\odot[\mathbf{x}_{n},\underbrace{{\bf g}(({\bf A}^{\top}{\bf d})^{(e)})}_{\begin{subarray}{c}\text{scaled cond.}\\ \text{features}\end{subarray}}]+\underbrace{\bm{\beta}_{t}}_{\begin{subarray}{c}\text{time}\\ \text{shift}\end{subarray}}\right),$ (18)
[R1-C2]
A sample $\bf m$} is drawn from the simulator,[R1-C3] together with optional conditional observations from Equation (14) and their adjoint ${\bf A}^{\top}{\bf d}$. Both tensors are passed through an encoder into a continuous space for flow matching. A stochastically sampled pair ${\bf z},t$ is drawn[R1-C3] and Equation (6) is used to generate the interpolated state ${\bf m}_{t}^{(e)}$ which serves as the training input for the [R1-C2] velocity network $\hat{\bf v}$, implemented as a 3D U-Net. The time embedding conditions the network on $t$, [R1-C2] while for CFM the conditional observations $\bf d$} are included via concatenated embeddings at each U-Net scale. Attention layers broadly distribute information spatially, while the skip connections transmit across resolutions.
The true velocity target $\mathbf{v}_{t}$ is computed using Equation (7) from the withheld pair $(\mathbf{m}^{(e)},\mathbf{z})$  (7)[R1-C3], and Equation (16) yields the reconstructed observations $\hat{\mathbf{d}}$. The predicted velocity $\hat{\mathbf{v}}_{t}^{(e)}$ is compared with $\mathbf{v}_{t}$ to form the conditional loss $\mathcal{L}_{\text{cond}}$, which backpropagates to update both the learnable time embeddings and the U-Net parameters.
Figure 3: Overview of our conditional flow-matching architecture. The diagram is inspired by the Latent Diffusion Models paper (Rombach et al., 2022).
[…]
3 Results
3.1 Unconditional Generation of Earth Models
@@
In our first set of experiments, we show that a flow matching model trained with suitable synthetic data in a continuously embedded space can generate realistic earth models. Unconditional generation is carried out using the learned velocity flow field and a numerical ODE solver [torchdiffeq]. After integration, the resulting sample is decoded into the categorical space, as formulated in (12), giving 3D categorical samples as shown in Figure 4.
@@
Figure 4: Three examples of unconditional earth models. Panels (a–c) $64\times 64\times 64$ resolution generative samples from $\pi_{{\bf m}}$ using flow matching. Each model represents a $3.84\,\text{km}\,\times 3.84\,\text{km}\,\times 3.84\,\text{km}$ volume of the Earth’s crust. Section cutaways into the generated model are added to show portions of the interior.
@@
These models are diverse, encompassing variations in topography, layer thickness and geometry, as well as dykes and intrusions. Such samples can support data generation for applications like seismic simulation [woollam2021seisbenchtoolboxmachine], geological segmentation and mapping [julka2023knowledgedistillationsegmentsam], and classification [Zhang2018GeoSeg].
3.2 Conditional Generation of Earth Models
@@
The standard approach to this problem is kriging [stein1999interpolation] or, more recently, interval kriging [Song2024Kriging], which estimates lithological probabilities between boreholes by accounting for vertical and horizontal variability in rock types. However, these methods can struggle with sparse data (few boreholes) and have the tendency to smooth interfaces when applied to categorical data. We demonstrate an alternative reconstruction method using a conditional flow matching model to interpolate between sparse boreholes. Because the technique is inherently generative, we can produce many different models that are consistent with the observed data and also with the underlying probability density function of the synthetic training dataset.
@@
Figure 6: Generative trajectory from Gaussian noise to geology conditioned on sparse data. Decoded categorical models at intermediary times $t\in[0,1]$ illustrate the continuous flow toward $\pi_{{\bf m}|{\bf d}}$.
@@
Figure 7: Three geological models from $\pi_{{\bf m}|{\bf d}}$ via numerical integration conditioned on available lithological data (top row) and their corresponding cut-section views (bottom row). The models offer a complete 3D $(64\times 64\times 64)$ resolution reconstruction, with the exterior surface of the models shown here for comparison. Cut-away views indicate changes throughout the internal range of the model, as reconstructions sample from a complex $\pi_{{\bf m}|{\bf d}}$ distribution.
@@
3.3 Conditional Feature Estimation and Reconstruction
Individual features of a reconstructed model, such as dike formations, can be analyzed in greater detail using an ensemble of samples from the conditional distribution $\pi_{\bf m}$. The planar dikes in the original model, the sparse dike data, and several reconstructions of their form are shown in Figure 8. The realizations show a high degree of variability due to the very limited conditional data.
Figure 8: Conditioned dike realizations. All models are rendered with a transparent exterior to emphasize only the 3D lithology of the planar dikes. (a) True dike geometry in situ (yellow and orange) with sparse surface/borehole data. (b) Twelve conditional realizations sampled from the model. The reconstructions demonstrate the diverse sampling possible from a learned $\pi_{{\bf m}|{\bf d}}$ distribution.
A large ensemble of samples can be used as a probability estimate for the likelihood of any individual voxel being part of the larger dike structure. A pattern emerges when plotting the extent of the planar dike based on the contour shape of probabilistic confidence intervals. Shown in Figure 9, a lower confidence interval corresponds to a larger and more aggressive extrapolation range from the existing boreholes. The smaller dike shown in the model has insufficient conditional data to produce an accurate reconstruction, a problem exacerbated by the true extent of the dike being largely obscured by sediment, see Figure 5 (a) and Figure 9 (b) for a comparison of the small surface signature to the larger subsurface extent.